A $2500 computer system can be leased for $79 per momth.After 3 years it can be purchased for $750. This is also the salvage value if the system was purchased originally. what is the effective annual rate for leasing the computer?
Answer #1

Monthly interest rate = r

PV = $ 2500 , monthly lease payment = $ 79

Time period = 3 years or 36 months

Salvage value = $ 750

2500 = 79/(1 + r) + 79/(1 + r)2 + .... + 79/(1 + r)36 + 750/(1 + r)36

r = 1.75%

Effective annual rate = (1 + 1.75%)12 - 1 = 23.1%